OpenStudy (anonymous):

How do I find the x and y intercepts of y = x(x - 4)

hartnn (hartnn):

to find x-intercept,put y=0, to find y-intercept ,put x=0 in the equation y = x(x - 4)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

So y = -4?

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

y = x(x - 4) Y Intercept: [0,b] y = 0(0-4) y = 0(0)+(0(-4)) y = 0 + 0 y = 0 y intercept: (0,0) --- X Intercept: [a,0] 0 = x(x-4)* 1.) x=0 2.) (x-4)=0 x = 4 x intercept/s: (4,0) and (0,0) Note. Use the zero factor/product property to solve the above: *Zero factor property: If ab=0, then either a=0 or b=0* --- Notes. The Fundamental Theorem of Algebra: If p(x) is a polynomial of degree n then p(x) will have exactly n zeroes, some of which may repeat. The Factor Theorem: "A polynomial equation f(x) has a factor (x-k) if and only if f(k) = 0." Quadratic Equations: *Zero factor property: If ab=0, then either a=0 or b=0* Factor Theorem of Algebra: http://en.wikipedia.org/wiki/Root-finding_algorithm Root-Finding Algorithm: http://en.wikipedia.org/wiki/Root-finding_algorithm Quadratic Equations and Zero Factor Property: http://tutorial.math.lamar.edu/Classes/Alg/SolveQuadraticEqnsI.aspx